OpenStudy (simplymarie_x):

The world’s largest round pizza (which happened to be a gluten-free pizza) was set in December 2012 by five Italian chefs. The diameter of this pizza was 131 feet. If this pizza was cut into 50 slices, what would be the area of each slice and the length of the crust for the slice? What is the difference between finding the area of a sector and the arc length? How do you find both?

OpenStudy (mathstudent55):

First, we need to assume that the 50 slices are all sectors of the circle and have the same central angle.

OpenStudy (mathstudent55):

Do you know the formula for the circumference of a circle?

OpenStudy (simplymarie_x):

I've done this so far: A = pi * r^2 pi = 3.14 r = d / 2 r = 131 / 2 = 65.5 Area = 3.14 * 65.5^2 Area of the whole pizza = 13471 feet^2 Area of one slice = 13471 / 50 = 269 square feet. Circumference = 2*pi * r = pi * d Circumference = 2 * 3.14 * 65.5 = 411 feet (rounded.) So 1/50 of the circumference = 411/50 = 8.22 feet. Is that right?

OpenStudy (mathstudent55):

Correct.

OpenStudy (simplymarie_x):

What would be the answers for "What is the difference between finding the area of a sector and the arc length? How do you find both?" I don't get that.

OpenStudy (mathstudent55):

The area of a sector is a fraction of the area of the circle, and the length of the arc of a sector is a fraction of the circumference of the circle.

OpenStudy (mathstudent55):

area of sector of circle is related to area of circle length of arc of sector is related to circumference of circle

OpenStudy (simplymarie_x):

Thanks so so much!

OpenStudy (mathstudent55):

You're welcome.